(Straight Line Depreciation) A new machine cost $100,000, and will increase inventory by $10,000, A/R by $15,000, and A/P by $5,000. This machine will reduce our expenses by $18,000 per year. The machine will be depreciated to a zero book value in 10 years, but could be sold for $12,000 at the end of 10 years. The marginal tax rate is 35%, and the cost of capital is 12%. Calculate the payback, NPV, IRR, and PI.
Answer #1

Payback

Initial Investment / annual payback

annual payback

(Asset cost – Residual Value) / Useful life of the asset

(100,000 - 15000 ) / 10 = 8500

So,

payback

100,000 / 8500 = 11.764

NPV

cashflow/(1+r)^t - initial investment

10000/(1+15000)^10 - 100,000

IRR

​IRR=NPV=t=1∑T​(1+r)tCt​​−C0​=0